# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 1.1.1 Cryptography in Ancient Egypt and Ancient Greece
Cryptography has its roots deep in ancient history, with various early civilizations employing basic cryptographic methods to protect their messages [1, 2, 3, 4, 5]. These techniques were essential for military, political, and diplomatic communications, where secrecy could mean the difference between success and failure.
In ancient Egypt, one of the earliest known examples of cryptography was found in the tombs of pharaohs [6]. Egyptian scribes used a method of encryption in which non-standard hieroglyphs were substituted for more common ones. This wasn’t designed for secrecy in the way we understand cryptography today but may have been used to obscure the message from the untrained.
The ancient Greeks took cryptography further, especially with the development of techniques like the Scytale. The Scytale was a tool used by the Spartans in military communication[7]. It involved wrapping a strip of parchment around a cylinder of a specific diameter. The message was written across the parchment, and when unwrapped, it would appear as random letters. Only by wrapping it around another cylinder of the same diameter could the original message be deciphered, making it one of the earliest examples of transposition ciphers.
Another early technique was Steganography, the practice of hiding a message within another, more benign message or medium[8, 9]. One of the famous examples is from Herodotus, who tells of a secret message being written underneath the wax of a wax tablet.
@@ 1.1.2 Caesar Cipher and Classical Encryption
One of the most famous classical encryption techniques is the Caesar Cipher [10, 11], named after Julius Caesar, who used this method to protect his military communications. The Caesar cipher is a type of substitution cipher where each letter in the plaintext is shifted by a fixed number of positions down the alphabet. For example, with a shift of 3, A would become D, B would become E, and so on.
@@ 1.2.1 Shannon’s Communication Theory
In 1949, Shannon published a seminal paper titled ”Communication Theory of Secrecy Systems[12]in which he applied the concepts of information theory to cryptography. Shannon introduced several key concepts that are still vital to modern cryptographic algorithms, such as entropy[13], confusion[14], and diffusion[14]:
  • Entropy: In cryptography, entropy refers to the uncertainty or randomness in a system[13]. The higher the entropy, the more secure a cryptographic system is. Shannon demonstrated that systems with low entropy are easier to predict and therefore more vulnerable to attacks.
  • Confusion: Confusion is a technique used to make the relationship between the key and the ciphertext as complex as possible[14]. This is done to prevent an attacker from finding patterns that could lead to determining the key. Substitution techniques, such as those used in substitution ciphers, are a way to introduce confusion.
  • Diffusion: Diffusion spreads the influence of each plaintext digit over many ciphertext digits, making it harder to deduce the plaintext by examining the ciphertext[14]. Transposition ciphers are an example of diffusion in action.
Shannon’s work laid the groundwork for understanding the fundamental properties that any secure encryption system should have. His theories directly influence the design of modern encryption algorithms, such as AES (Advanced Encryption Standard)[15] and RSA (Rivest–Shamir–Adleman)[16], which employ high levels of confusion and diffusion to ensure security.
@@ 1.2.2 Security Definitions in Cryptography
In cryptography, the goals of securing communication or data can be summarized using three core principles: Confidentiality, Integrity, and Authenticity. These are often referred to as the CIA triad[17].
@@
Cryptographic security models, such as Kerckhoffs’s Principle[18], emphasize that the security of a system should rely on the secrecy of the key rather than the secrecy of the algorithm itself. This means that cryptographic algorithms should be open to public scrutiny, but the key must remain secret.
In designing cryptographic algorithms, these principles guide how we approach problems like key management, authentication protocols, and secure communication channels. For instance, modern encryption methods like AES[19] use both confusion and diffusion to ensure confidentiality, while digital signatures ensure authenticity, and hash functions help maintain integrity.
@@ 1.3.1 Ensuring Data Confidentiality and Integrity
One of the key benefits of blockchain technology is the ability to ensure that data remains both confidential and secure from unauthorized changes[20]. Cryptographic techniques, specifically symmetric and asymmetric encryption, are employed to safeguard sensitive information.
@@
Symmetric encryption involves using the same key for both encryption and decryption[21]. This method is highly efficient and suitable for encrypting large amounts of data, making it ideal for protecting user data within a blockchain network. However, the challenge with symmetric encryption is securely sharing the secret key, as anyone who has the key can decrypt the data.
@@
Asymmetric encryption, also known as public-key cryptography, uses a pair of keys—a public key and a private key[22]. The public key is shared openly and used for encryption, while the private key is kept secret and used for decryption. This method provides a higher level of security compared to symmetric encryption and is commonly used in blockchain systems to protect users’ identities and ensure that only authorized parties can access specific information.
@@
While encryption protects the confidentiality of data, cryptography also ensures that the data remains unaltered. Integrity in blockchain is often achieved through cryptographic hash functions[23], which we will cover in more detail in a later section. For now, consider that any modification to a block in the blockchain would change the hash of the block, making tampering evident.
@@ 1.3.2 Achieving Trustlessness: Cryptography as the Foundation
The concept of trustlessness refers to the elimination of the need for trusted third parties (intermediaries) in a transaction[24]. Blockchain achieves this through cryptography, where the rules of the system and trust are embedded in mathematical principles, not reliant on human intervention[23].
@@
In blockchain, public key cryptography is the foundation of user identity[25]. Each participant in the network has a cryptographic key pair—a public key that can be shared and a private key that is kept secret.
@@
Digital signatures are another crucial cryptographic tool used in blockchain to ensure trustlessness[26]. When a user signs a transaction with their private key, it generates a digital signature that can be verified using their public key. This mechanism allows the blockchain to verify the authenticity of transactions without intermediaries.
@@ 1.3.3 The Different Roles of Encryption and Hashing in Blockchain
Encryption, as discussed earlier, is primarily used to ensure that data remains confidential and can only be accessed by authorized parties[27]. In blockchain, encryption is mainly employed to protect user data and communication between nodes.
@@
In contrast to encryption, hashing does not aim to protect confidentiality but rather ensures data integrity[28]. A hash function takes an input (or ”message”) and returns a fixed-size string of bytes[29]. Even a small change in the input drastically changes the output hash, making it easy to detect tampering.
@@
  • Block identification: Each block in a blockchain contains a hash of the previous block, creating a chain of blocks that ensures any alteration would break the chain[30].
  • Proof of Work (PoW): In Bitcoin, miners compete to solve a cryptographic puzzle, which involves finding a hash value below a certain threshold[31]. This process secures the network and prevents double-spending.
  • Transaction integrity: Each transaction is hashed and included in a block, making it immutable once added to the blockchain[32].
@@
Proof of Work is a consensus algorithm that relies on hashing to secure the blockchain[31]. In PoW-based blockchains like Bitcoin, miners compete to solve a complex mathematical problem that involves finding a hash value. Once a miner finds the correct hash, they can add the new block to the blockchain, and the process starts again.
@@
13 return guess_hash[:4] == "$0$"
@@ 2.1 Definition and Characteristics of Symmetric Encryption
Symmetric encryption is a cryptographic method in which the same key is used for both encryption and decryption[33]. It is one of the most fundamental and widely used encryption techniques due to its simplicity and efficiency. In this type of encryption, the sender encrypts the plaintext using a key, and the receiver decrypts the ciphertext with the same key[34]. This method ensures confidentiality, as long as the key remains secret between the communicating parties.
The main characteristics of symmetric encryption include[35]:
@@ 2.1.1 The Key Sharing Problem
One of the major challenges in symmetric encryption is the secure distribution of the encryption key[36]. Since the same key is used for both encryption and decryption, it must be shared securely between the communicating parties. This is known as the key sharing problem.
If the key is intercepted during transmission, the security of the encrypted data is compromised. Therefore, securely exchanging the encryption key between the sender and receiver is a critical issue.
@@
  • Key Exchange Protocols: Protocols such as the Diffie-Hellman key exchange[37] allow two parties to securely share a secret key over an unsecured communication channel without revealing the key to any third party.
@@
  • Hybrid Encryption: A common approach is to use asymmetric encryption to securely exchange the symmetric key[38]. For example, in SSL/TLS protocols[39], asymmetric encryption is used to exchange a session key, which is then used for faster symmetric encryption during the communication.
@@ 2.1.2 Security Analysis of Symmetric Encryption
  • Efficiency: Symmetric encryption is computationally efficient, making it ideal for encrypting large datasets or streaming data[40].
  • Simplicity: The implementation of symmetric encryption algorithms is generally more straightforward than asymmetric encryption[41].
  • Low Computational Cost: Symmetric encryption is less resource-intensive compared to its asymmetric counterpart, which makes it suitable for devices with limited computational power[42].
@@
  • Brute-force Attacks: Symmetric encryption is vulnerable to brute-force attacks, where an attacker systematically attempts every possible key[43]. To mitigate this, modern symmetric algorithms use sufficiently long key sizes (e.g., AES with a 256-bit key)[44].
  • Key Distribution: As mentioned earlier, securely distributing the key remains a critical challenge[36].
  • Key Management: Managing and securely storing keys, especially in large-scale systems with multiple parties, can be difficult. If keys are not securely stored, the encryption becomes vulnerable to key exposure[45].
@@ 2.2 Stream Ciphers
Stream ciphers are a type of symmetric encryption that encrypt data one bit or byte at a time, as opposed to block ciphers, which encrypt data in fixed-size blocks[46]. This allows stream ciphers to be highly efficient and low latency, making them ideal for applications like real-time communications.
@@ 2.2.1 Encryption and Decryption Process of Stream Ciphers
The encryption process in stream ciphers involves generating a keystream, a pseudo-random sequence of bits, which is XORed with the plaintext to produce ciphertext[47]. Similarly, the decryption process XORs the ciphertext with the same keystream to retrieve the original plaintext.
@@ 2.2.2 The RC4 Algorithm and Its Applications
RC4 is one of the most well-known stream ciphers[48]. It was widely used in protocols such as WEP (Wired Equivalent Privacy)[49] and SSL/TLS[50]. RC4 operates by initializing a key-scheduling algorithm (KSA)[51] to create a keystream from a secret key. This keystream is XORed with the plaintext to produce ciphertext.
However, RC4 has been found to have several vulnerabilities, particularly due to biases in the keystream[52]. Attackers can exploit these biases to recover parts of the key or plaintext, making RC4 insecure for many applications.
@@
  • Pseudo-random Generation Algorithm (PRGA): The PRGA generates the keystream, which is XORed with the plaintext[53].
@@ 2.2.3 Advantages and Limitations of Stream Ciphers
  • Low Latency: Since stream ciphers encrypt data one bit or byte at a time, they are well-suited for environments where low latency is essential, such as real-time communication[54].
  • Efficiency: Stream ciphers are typically faster than block ciphers for encrypting small amounts of data or streaming data[55].
  • Simplicity in Hardware: Stream ciphers can be implemented efficiently in hardware, making them suitable for embedded systems and constrained environments[56].
@@
  • Key Reuse Vulnerability: One of the biggest weaknesses of stream ciphers is the susceptibility to attacks if the same key is used more than once. Reusing the same key with different data can lead to key recovery attacks[57].
  • Known Attacks: Stream ciphers like RC4 have known vulnerabilities[52]. For example, biases in the RC4 keystream can be exploited to break its encryption.
  • Synchronization Requirements: In some stream cipher implementations, both sender and receiver must remain synchronized for the decryption process to work correctly[58]. Loss of synchronization can result in data corruption.

### v3

@@
   Xinyuan Song
Emory University
xinyuan.song@emory.edu
@@ 1.1.1 Cryptography in Ancient Egypt and Ancient Greece
Cryptography has its roots deep in ancient history, with various early civilizations employing basic cryptographic methods to protect their messages [subramani2023review, aumasson2024serious, thabit2023cryptography, pillai2024analyzing, zulkifli2007evolution]. These techniques were essential for military, political, and diplomatic communications, where secrecy could mean the difference between success and failure.
In ancient Egypt, one of the earliest known examples of cryptography was found in the tombs of pharaohs [polis2023guide]. Egyptian scribes used a method of encryption in which non-standard hieroglyphs were substituted for more common ones. This wasn’t designed for secrecy in the way we understand cryptography today but may have been used to obscure the message from the untrained.
The ancient Greeks took cryptography further, especially with the development of techniques like the Scytale. The Scytale was a tool used by the Spartans in military communication[guerra2024embryonic]. It involved wrapping a strip of parchment around a cylinder of a specific diameter. The message was written across the parchment, and when unwrapped, it would appear as random letters. Only by wrapping it around another cylinder of the same diameter could the original message be deciphered, making it one of the earliest examples of transposition ciphers.
Another early technique was Steganography, the practice of hiding a message within another, more benign message or medium[rustad2023digital, jamil1999steganography]. One of the famous examples is from Herodotus, who tells of a secret message being written underneath the wax of a wax tablet.
@@ 1.1.2 Caesar Cipher and Classical Encryption
One of the most famous classical encryption techniques is the Caesar Cipher [ardiansyah2023implementasi, noviyanti2analysis], named after Julius Caesar, who used this method to protect his military communications. The Caesar cipher is a type of substitution cipher where each letter in the plaintext is shifted by a fixed number of positions down the alphabet. For example, with a shift of 3, A would become D, B would become E, and so on.
@@ 1.2.1 Shannon’s Communication Theory
In 1949, Shannon published a seminal paper titled ”Communication Theory of Secrecy Systems[shannon1949communication]in which he applied the concepts of information theory to cryptography. Shannon introduced several key concepts that are still vital to modern cryptographic algorithms, such as entropy[simion2020entropy], confusion[hu2021quantum], and diffusion[hu2021quantum]:
  • Entropy: In cryptography, entropy refers to the uncertainty or randomness in a system[simion2020entropy]. The higher the entropy, the more secure a cryptographic system is. Shannon demonstrated that systems with low entropy are easier to predict and therefore more vulnerable to attacks.
  • Confusion: Confusion is a technique used to make the relationship between the key and the ciphertext as complex as possible[hu2021quantum]. This is done to prevent an attacker from finding patterns that could lead to determining the key. Substitution techniques, such as those used in substitution ciphers, are a way to introduce confusion.
  • Diffusion: Diffusion spreads the influence of each plaintext digit over many ciphertext digits, making it harder to deduce the plaintext by examining the ciphertext[hu2021quantum]. Transposition ciphers are an example of diffusion in action.
Shannon’s work laid the groundwork for understanding the fundamental properties that any secure encryption system should have. His theories directly influence the design of modern encryption algorithms, such as AES (Advanced Encryption Standard)[smid2021development] and RSA (Rivest–Shamir–Adleman)[cohen2011architecture], which employ high levels of confusion and diffusion to ensure security.
@@ 1.2.2 Security Definitions in Cryptography
In cryptography, the goals of securing communication or data can be summarized using three core principles: Confidentiality, Integrity, and Authenticity. These are often referred to as the CIA triad[khansa2014assessing].
@@
Cryptographic security models, such as Kerckhoffs’s Principle[ke2019generative], emphasize that the security of a system should rely on the secrecy of the key rather than the secrecy of the algorithm itself. This means that cryptographic algorithms should be open to public scrutiny, but the key must remain secret.
In designing cryptographic algorithms, these principles guide how we approach problems like key management, authentication protocols, and secure communication channels. For instance, modern encryption methods like AES[inam2023new] use both confusion and diffusion to ensure confidentiality, while digital signatures ensure authenticity, and hash functions help maintain integrity.
@@ 1.3.1 Ensuring Data Confidentiality and Integrity
One of the key benefits of blockchain technology is the ability to ensure that data remains both confidential and secure from unauthorized changes[lin2024symmetry]. Cryptographic techniques, specifically symmetric and asymmetric encryption, are employed to safeguard sensitive information.
@@
Symmetric encryption involves using the same key for both encryption and decryption[sudeva2017symmetric]. This method is highly efficient and suitable for encrypting large amounts of data, making it ideal for protecting user data within a blockchain network. However, the challenge with symmetric encryption is securely sharing the secret key, as anyone who has the key can decrypt the data.
@@
Asymmetric encryption, also known as public-key cryptography, uses a pair of keys—a public key and a private key[mohamad2021research]. The public key is shared openly and used for encryption, while the private key is kept secret and used for decryption. This method provides a higher level of security compared to symmetric encryption and is commonly used in blockchain systems to protect users’ identities and ensure that only authorized parties can access specific information.
@@
While encryption protects the confidentiality of data, cryptography also ensures that the data remains unaltered. Integrity in blockchain is often achieved through cryptographic hash functions[kuznetsov2021performance], which we will cover in more detail in a later section. For now, consider that any modification to a block in the blockchain would change the hash of the block, making tampering evident.
@@ 1.3.2 Achieving Trustlessness: Cryptography as the Foundation
The concept of trustlessness refers to the elimination of the need for trusted third parties (intermediaries) in a transaction[bruun2020infrastructures]. Blockchain achieves this through cryptography, where the rules of the system and trust are embedded in mathematical principles, not reliant on human intervention[kuznetsov2021performance].
@@
In blockchain, public key cryptography is the foundation of user identity[tewari2019blockchain]. Each participant in the network has a cryptographic key pair—a public key that can be shared and a private key that is kept secret.
@@
Digital signatures are another crucial cryptographic tool used in blockchain to ensure trustlessness[fang2020digital]. When a user signs a transaction with their private key, it generates a digital signature that can be verified using their public key. This mechanism allows the blockchain to verify the authenticity of transactions without intermediaries.
@@ 1.3.3 The Different Roles of Encryption and Hashing in Blockchain
Encryption, as discussed earlier, is primarily used to ensure that data remains confidential and can only be accessed by authorized parties[zhang2018ensuring]. In blockchain, encryption is mainly employed to protect user data and communication between nodes.
@@
In contrast to encryption, hashing does not aim to protect confidentiality but rather ensures data integrity[mohammed2020maintaining]. A hash function takes an input (or ”message”) and returns a fixed-size string of bytes[pappu2002physical]. Even a small change in the input drastically changes the output hash, making it easy to detect tampering.
@@
  • Block identification: Each block in a blockchain contains a hash of the previous block, creating a chain of blocks that ensures any alteration would break the chain[monrat2019survey].
  • Proof of Work (PoW): In Bitcoin, miners compete to solve a cryptographic puzzle, which involves finding a hash value below a certain threshold[cojocaru2023quantum]. This process secures the network and prevents double-spending.
  • Transaction integrity: Each transaction is hashed and included in a block, making it immutable once added to the blockchain[patel2021review].
@@
Proof of Work is a consensus algorithm that relies on hashing to secure the blockchain[cojocaru2023quantum]. In PoW-based blockchains like Bitcoin, miners compete to solve a complex mathematical problem that involves finding a hash value. Once a miner finds the correct hash, they can add the new block to the blockchain, and the process starts again.
@@
13 return guess_hash[:4] == "0000"
@@ 2.1 Definition and Characteristics of Symmetric Encryption
Symmetric encryption is a cryptographic method in which the same key is used for both encryption and decryption[Nielson2019SymmetricET]. It is one of the most fundamental and widely used encryption techniques due to its simplicity and efficiency. In this type of encryption, the sender encrypts the plaintext using a key, and the receiver decrypts the ciphertext with the same key[Hellwig2020BlockchainCP]. This method ensures confidentiality, as long as the key remains secret between the communicating parties.
The main characteristics of symmetric encryption include[Gharat2014OverviewOS]:
@@ 2.1.1 The Key Sharing Problem
One of the major challenges in symmetric encryption is the secure distribution of the encryption key[yashaswini2015key]. Since the same key is used for both encryption and decryption, it must be shared securely between the communicating parties. This is known as the key sharing problem.
If the key is intercepted during transmission, the security of the encrypted data is compromised. Therefore, securely exchanging the encryption key between the sender and receiver is a critical issue[].
@@
  • Key Exchange Protocols: Protocols such as the Diffie-Hellman key exchange[Canetti2001AnalysisOK] allow two parties to securely share a secret key over an unsecured communication channel without revealing the key to any third party.
@@
  • Hybrid Encryption: A common approach is to use asymmetric encryption to securely exchange the symmetric key[Simmons1979SymmetricAA]. For example, in SSL/TLS protocols[2021RSAKS], asymmetric encryption is used to exchange a session key, which is then used for faster symmetric encryption during the communication.
@@ 2.1.2 Security Analysis of Symmetric Encryption
  • Efficiency: Symmetric encryption is computationally efficient, making it ideal for encrypting large datasets or streaming data[Raigoza2016EvaluatingPO].
  • Simplicity: The implementation of symmetric encryption algorithms is generally more straightforward than asymmetric encryption[AlShabi2019ASO].
  • Low Computational Cost: Symmetric encryption is less resource-intensive compared to its asymmetric counterpart, which makes it suitable for devices with limited computational power[nour2024securing].
@@
  • Brute-force Attacks: Symmetric encryption is vulnerable to brute-force attacks, where an attacker systematically attempts every possible key[Marinakis2013MinimumKL]. To mitigate this, modern symmetric algorithms use sufficiently long key sizes (e.g., AES with a 256-bit key)[Blaze1996MinimalKL].
  • Key Distribution: As mentioned earlier, securely distributing the key remains a critical challenge[yashaswini2015key].
  • Key Management: Managing and securely storing keys, especially in large-scale systems with multiple parties, can be difficult. If keys are not securely stored, the encryption becomes vulnerable to key exposure[Damera2012SecureSO].
@@ 2.2 Stream Ciphers
Stream ciphers are a type of symmetric encryption that encrypt data one bit or byte at a time, as opposed to block ciphers, which encrypt data in fixed-size blocks[cryptoeprint:2004/094]. This allows stream ciphers to be highly efficient and low latency, making them ideal for applications like real-time communications.
@@ 2.2.1 Encryption and Decryption Process of Stream Ciphers
The encryption process in stream ciphers involves generating a keystream, a pseudo-random sequence of bits, which is XORed with the plaintext to produce ciphertext[Gorbenko2017TheRO]. Similarly, the decryption process XORs the ciphertext with the same keystream to retrieve the original plaintext.
@@ 2.2.2 The RC4 Algorithm and Its Applications
RC4 is one of the most well-known stream ciphers[Jindal2015ASO]. It was widely used in protocols such as WEP (Wired Equivalent Privacy)[Stoic2012RC4SC] and SSL/TLS[Khine2009ANV]. RC4 operates by initializing a key-scheduling algorithm (KSA)[ohigashi2008new] to create a keystream from a secret key. This keystream is XORed with the plaintext to produce ciphertext.
However, RC4 has been found to have several vulnerabilities, particularly due to biases in the keystream[Paterson2014BigBH]. Attackers can exploit these biases to recover parts of the key or plaintext, making RC4 insecure for many applications.
@@
  • Pseudo-random Generation Algorithm (PRGA): The PRGA generates the keystream, which is XORed with the plaintext[Maity2017AnEL].
@@ 2.2.3 Advantages and Limitations of Stream Ciphers
  • Low Latency: Since stream ciphers encrypt data one bit or byte at a time, they are well-suited for environments where low latency is essential, such as real-time communication[Venkatesulu2016ASC].
  • Efficiency: Stream ciphers are typically faster than block ciphers for encrypting small amounts of data or streaming data[Sharif2010PerformanceAO].
  • Simplicity in Hardware: Stream ciphers can be implemented efficiently in hardware, making them suitable for embedded systems and constrained environments[Llise2005EfficientIO].
@@
  • Key Reuse Vulnerability: One of the biggest weaknesses of stream ciphers is the susceptibility to attacks if the same key is used more than once. Reusing the same key with different data can lead to key recovery attacks[Stone2020RethinkingTW].
  • Known Attacks: Stream ciphers like RC4 have known vulnerabilities[Paterson2014BigBH]. For example, biases in the RC4 keystream can be exploited to break its encryption.
  • Synchronization Requirements: In some stream cipher implementations, both sender and receiver must remain synchronized for the decryption process to work correctly[Armknecht2004ExtendingTR]. Loss of synchronization can result in data corruption.